# From Text to [Uncaptioned image]: How PEFT-Driven Personality Manipulation Unleashes the Emoji Potential in LLMs

### v1

@@ A.8 Manipulation Validation
Here, the model was asked to generate the top 5 tokens that best matched the personality from the generated text. Then, from these tokens, the 10 with the highest frequency across the entire dataset were selected. Figures 16-20 show the results obtained from this analysis.
From Figures 16-20, it is evident that both Llama-2-7B-Chat and Mistral-7B-Instruct utilise emojis with intention, rather than as random outputs. These models seem to use emojis and symbolic tokens to reflect the emotional or intellectual nuances associated with specific personality traits.
@@
Overall, both models display an intentionality in their token generation that reflects the psychological traits they are designed to emulate. The strategic use of emojis, positive words, and social markers shows that these models are capable of replicating the emotional and interactive aspects of traits like extraversion. This shows that AI models are becoming more advanced, as they are better able to reflect complex human behaviors and emotions, making them more similar to how humans think and interact.

### v2

@@ A.8 Manipulation Validation and ICL Explainability
Here, the model was asked to generate the top 5 tokens that best matched the personality from the generated text. Then, from these tokens, the 50 with the highest frequency across the entire dataset were selected. Figures 18-22 show the results obtained from this analysis. These figures show only top 10 tokens due to space issues.
From Figures 18-22, it is evident that both Llama-2-7B-Chat and Mistral-7B-Instruct utilise emojis with intention, rather than as random outputs. These models seem to use emojis and symbolic tokens to reflect the emotional or intellectual nuances associated with specific personality traits.
@@
Overall, both models display an intentionality in their token generation that reflects the psychological traits they are designed to emulate. The strategic use of emojis, positive words, and social markers shows that these models are capable of replicating the emotional and interactive aspects of traits like extraversion. This shows that AI models are becoming more advanced, as they are better able to reflect complex human behaviours and emotions, making them more similar to how humans think and interact.